# CO3: Contrasting Concepts Compose Better

### v1

@@ Abstract
We propose to improve multi‑concept prompt fidelity in text‑to‑image diffusion models. We begin with common failure cases—prompts like “a cat and a dog” that sometimes yields images where one concept is missing, faint, or colliding awkwardly with another. We hypothesize that this happens when the diffusion model drifts into mixed modes that over‑emphasize a single concept it learned strongly during training. Instead of re-training, we introduce a corrective sampling strategy that steers away from regions where the joint prompt behavior overlaps too strongly with any single concept in the prompt. The goal is to steer towards “pure” joint modes where all concepts can coexist with balanced visual presence.. We further show that existing multi‑concept guidance schemes can operate in unstable weight regimes that amplify imbalance; we characterize favorable regions and adapt sampling to remain within them. Our approach, CO3, is plug‑and‑play, requires no model tuning, and complements standard classifier‑free guidance. Experiments on diverse multi-concept prompts indicate improvements in concept coverage, balance and robustness, with fewer dropped or distorted concepts compared to standard baselines and prior compositional methods. Results suggest that lightweight corrective guidance can substantially mitigate brittle semantic alignment behavior in modern diffusion systems.
@@ 1 Introduction
Recent diffusion models (Ho et al., 2020; Ramesh et al., 2022a; Rombach et al., 2022) have ushered significant breakthroughs in Text-to-Image (T2I) synthesis, producing high-fidelity images from textual descriptions. However, ensuring the generated images faithfully adhere to the prompt, a challenge known as semantic alignment (Chefer et al., 2023; taihang Hu et al., 2024; Liu et al., 2023), remains a problem. Concretely, for a given prompt $C$, T2I models like StableDiffusion (Rombach et al., 2022) sample from the modes (or high probability regions) of the learned distribution, $p(x|C)$. While such models can produce high resolution images in general, every so often, the results are surprisingly misaligned even for very simple prompts containing few concepts, e.g., CC=“a cat and a dog”. Diagnosing exactly why this behavior emerges periodically is difficult. It is conceivable that the complex training process in high dimensions, especially in conjunction with text embeddings, creates some problematic modes in $p(x|C)$.
We hypothesize that problematic modes in $p(x|C)$ arise when they overlap with modes of individual concept distribution $p(x|c_{i})$. Such an overlap biases the generation toward a single concept $c_{i}$, reducing the prominence of others. For instance, across images of $c_{1}=\text{``cat''}$ in the training dataset, a few may have an inconspicuous or partial $c_{2}=\text{``dog''}$ in the background. This image may still fall under the mode of $p(x|C)$. We attribute this to training instabilities and relatively less coverage of multi-concept prompts $C$, which cause the model to assign high probability even to weakly conforming images. Said differently, an image of a big cat and an inconspicuous dog can get assigned high probabilities under $p(x|C)$, causing semantic misalignment.
@@
Figure 1 illustrates the intuition behind our proposal. Our corrector distribution $\tilde{p}(x|C)$ assigns low probability to regions where $p(x|C)$ overlaps with individual $p(x|c_{i})$; we deem them as degenerate modes dominated by a single concept. By suppressing these overlaps, the corrector emphasizes pure $p(x|C)$ modes where all concepts coexist without one overwhelming the others. From a probabilistic perspective, this acts as a corrective factor: while $p(x|C)$ may assign high probability to weakly conforming images due to training noise or limited multi-concept data, dividing by the marginals removes this bias and sharpens the distribution toward genuine multi-concept samples. As a result, the modes we target are more semantically aligned and less prone to concept suppression or distortion.
Correction sampling from $\tilde{p}(x|C)$ can be achieved by composing scores from constituent component distributions $\nabla_{x}\log p(x|C),\{\nabla_{x}\log p(x|c_{i})\}_{i}$ (Liu et al., 2022). While there are many ways to compose, we analyse and show that composition through weighted sum of Tweedie-means—in the Tweedie denoised space—offers a more general framework that subsumes existing approaches.
In particular, we show that two classes of correction sampling—noise-resampling and latent correction (Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024; Kwon & Ye, 2025)—become special cases of Tweedie-mean composition under different weighting schemes. This allows us to design a hybrid composition framework that serves the purpose of resampling at time $T$, and then toggles to latent correction at later steps. By latent, we mean that the correction steps are accomplished in between each DDIM time step, allowing CO3 to be a plug-and-play, model-agnostic and gradient-free approach for T2I models. Comparison of CO3 against SOTA baselines shows stronger semantic alignment to prompts (measured using multiple metrics), giving empirical evidence of our hypothesis.
Figure 1: The figure illustrates our hypothesis on mode overlap using a simple 2D toy example. (a) Two modes of the distribution $p_{t}(x|\textit{"a cat and a dog"})$ (in green contour) has significant overlap with the modes of the individual concept distributions $p_{t}(x|\textit{"a cat"})$ (in red contour) and $p_{t}(x|\textit{"a dog"})$ (in orange contour). (b) The proposed corrector distribution $p_{t}(x|\texttt{"a cat and a dog"})/(p_{t}(x|\textit{"a cat"})p_{t}(x|\textit{"a dog"}))$ suppresses these overlaps, steering the generation away from problematic modes. The arrows indicate the denoising directions.
@@ 2 Background
In practice, most T2I models adopt classifier-free guidance (CFG) (Ho & Salimans, 2022), where predictions from both the conditional and unconditional models are combined:
@@
where $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]:=x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}^{\lambda,c}$ is the Tweedie mean from the CFG noise at $t$. Thus, CFG in the Tweedie framework can be interpreted as modifying the denoised estimate $\hat{x}_{0}$ to $\hat{x}_{0}^{\lambda,c}$ before the deterministic step to $x_{t-1}$ (Chung et al., 2024; Kwon & Ye, 2023).
@@
$\tilde{p}_{0}(x_{0}\mid c_{1},\ldots,c_{K})\;\propto\;\prod_{k=1}^{K}p_{0}(x_{0}\mid c_{k}).$ (9)
@@
3 CO3: Contrasting Concepts to Compose
We aim to combine the strengths of correction-based approaches and composable diffusion. On one hand, correction-based methods are powerful and explicitly improve compositional alignment, but they are subject to the complex gradient of the base model. On the other hand, composable diffusion is fully model-agnostic, but suffers from poor performance because linear score composition is not consistent with the diffusion forward process.
To take advantage of both, assume that prompt $C$ can be decomposed into constituent concepts $\{c_{1},c_{2},\ldots,c_{K}\}$. We propose an explicit Concept Contrasting corrector distribution based on our hypothesis on mode overlap discussed in Sec. 1. Specifically, we define the corrector distribution at each timestep $t$ as:
@@ 3 CO3: Contrasting Concepts to Compose
To sample from the unnormalized probability distribution in Eq. 12, an well-known approach is to compose the concept distributions $\{p_{t}(x_{t}\mid c_{k})\}_{k}$ in the space of score functions (Liu et al., 2022; Du et al., 2024). We break-away from this approach and compose the distribution in the Tweedie mean space; we show how this offers a more general framework for composition.
In Tweedie-denoised space we define composition as:
@@
8:  $x_{t}\leftarrow\text{DDIM\_step}(x_{t})$ $\triangleright$ Inverse DDIM
@@
A valid compositional Tweedie mean $\tilde{x}_{tweedie}$ in Eq. 13 must be one that can be expressed in the definition given in equation 5 as DDIM goes through a series of Tweedie means for image generation as described in section 2. Lemma 1 shows that this is guaranteed only when the weights satisfy the normalization condition $\sum_{i}w_{i}=1$.
Lemma 1.
Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
@@
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
@@ 4.1 Experimental Setup
Implementation details: We implement our method on SDXL (Podell et al., 2023) using a 50-step DDIM sampler with guidance scale $\lambda=5.0$. CO3 is applied during the first $20\%$ of denoising steps ($T_{c}=0.8T$), with the initial three steps reserved for resampling. For weight modulation, we set $\beta=0.8$. All experiments are performed on a single NVIDIA A6000 GPU (48GB). Additional hyperparameter details are provided in the Appendix section A.3.
Evaluation benchmark and metrics: We evaluate our method on the benchmark dataset from Chefer et al. (2023), which includes three categories of prompts: Animal–Animal, Animal–Object, and Object–Object. While the benchmark is limited to these cases, CO3 naturally extends to more complex prompts involving arbitrary numbers of concepts. For evaluation, we adopt BLIP-VQA (Huang et al., 2023) and ImageReward (Xu et al., 2023), both widely used in recent works (taihang Hu et al., 2024; Hu et al., 2024) to measure the faithfulness of generated images to input prompts. Notably, ImageReward is a learned reward model trained on human preferences, capturing aspects such as prompt adherence, aesthetics, and overall quality. For both metrics, higher scores indicate better alignment.
Comparison methods: We compare CO3 against state-of-the-art approaches, including optimization-based correction methods (Attend-and-Excite (Chefer et al., 2023), SynGen (Rassin et al., 2023), Divide-and-Bind (Liu et al., 2023), ToME (taihang Hu et al., 2024)), the composable generation method Composable Diffusion (Liu et al., 2022), and the noise-optimization method InitNO (Guo et al., 2024). All optimization-based approaches are architecture-dependent and require gradient computations at inference, with InitNO being model-agnostic but still incurring costly backpropagation. In contrast, only Composable Diffusion and our proposed CO3 are purely sampling-based, model-agnostic, and completely gradient-free.
@@ 4.2 Experimental Results
Properties BLIP-VQA ImageReward
training-free gradient-free model-agnostic animals animals-objects objects animals animals-objects objects
SD1.5(Rombach et al., 2022) - 0.3239 0.5958 0.2730 -0.2733 0.4262 -0.5521
Attend-excite(Chefer et al., 2023) 0.6980 0.7865 0.5155 0.8244 1.2380 0.8741
SynGenRassin et al. (2023) 0.3348 0.7689 0.6595 -0.2445 0.9838 0.7315
Divide & Bind 0.7201 0.8399 0.5887 0.8499 1.2516 0.8134
InitNO(Guo et al., 2024) 0.7264 0.7998 0.5406 1.0082 1.3927 1.1383
SDXLPodell et al. (2023) - 0.6950 0.8654 0.4926 0.7820 1.5574 0.6789
SynGen(SDXL)Rassin et al. (2023) 0.6816 0.8578 0.4652 0.6998 1.5622 0.6441
ToMe(taihang Hu et al., 2024) 0.6257 0.8808 0.6440 0.3895 1.5736 1.0118
Compose-diffusion(Liu et al., 2022) 0.2846 0.5656 0.4529 -1.1399 -0.2068 -0.0955
CO3(Ours) 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158
Table 1: Quantitative comparison of different methods on compositional generation tasks. We evaluate the generated images using two metrics: BLIP-VQA and ImageReward. Top performing model is highlighted in Black and 2nd best in Blue. Higher the score the better.
@@
Table 1 demonstrates the extent to which CO3, despite being model-agnostic and optimization-free, competes or outperforms optimization-based and model-agnostic baselines on ImageReward. Gains are substantive in the Animal–Animal and Animal–Object categories. On BLIP-VQA, CO3 delivers the best performance in these two categories and improves over the base SDXL on the Object category, which contains prompts of the form: “a [color1] [object1] and a [color2] [object2].”
Unlike methods such as ToMe taihang Hu et al. (2024), SynGen (Rassin et al., 2023), Attend-and-Excite (Chefer et al., 2023), and Divide-and-Bind Liu et al. (2023), which rely on model-specific architectures or explicit attention-based binding losses, CO3 is architecture-independent, sampling-based, and free of explicit subject–attribute binding. Interestingly, by simply targeting “pure” modes under $p(x\mid C)$, CO3 discovers layouts that better capture multi-object relations (Animals–Animals, Animal–Objects) and finds high-probability regions that preserve subject–attribute bindings (Objects). This performance gain empirically validates our hypothesis on mode overlap and also highlights CO3’s effectiveness in correcting the compositional limitations of SDXL.
Qualitative Comparison: Figure 3 compares CO3 with several baselines. Columns 1–2 show results from Divide-Bind (Liu et al., 2023) and Attend-and-Excite (Chefer et al., 2023) (based on SDv1.5 (Rombach et al., 2022)), while the remaining columns compare against Composable Diffusion (Liu et al., 2022), SynGen (Rassin et al., 2023), and ToMe (taihang Hu et al., 2024). We illustrate common issues in compositional alignment—concept missing, attribute mixing, and object binding—across all categories. Divide-Bind and Attend-and-Excite enhance the attention layer of the base model, but often yield cropped or incomplete concepts (rows 1, 3). Composable Diffusion suffers from missing or merged concepts (rows 2, 3). SynGen and ToMe perform well on Object prompts of the form “a [color1] [object1] and a [color2] [object2],” but still exhibit overlapping or mixed concepts (rows 2, 3; columns 5, 6). In contrast, CO3 achieves stronger object binding in the Animals and Animal–Objects categories while also improving results on Objects.
@@
Figure 4 extends qualitative results to complex prompts from the Complex category of Huang et al. (2023), comparing CO3 with the top-performing SDXL models (Table 1). The first row shows prompts involving spatial relations (e.g., “on,” “to the left of”), while the second row depicts more descriptive prompts with multiple interacting concepts. SynGen and ToMe successfully bind concepts to their subjects but often miss concepts or relations. By mitigating dominant-concept bias, CO3 remains effective even for complex prompts with multiple relations.
Figure 4: Qualitative comparison of CO3 with competing methods on complex prompts.
Ablations Studies are preformed on the key components of CO3, starting from the base SDXL model (Table 2). Adding resampling in the first three DDIM steps substantially improves average performance, particularly for the Animals and Animal–Objects categories. Since image layout is largely determined in the early timesteps, CO3-resampler iteratively explores better initial noise configurations that align with the prompt. Incorporating CO3-corrector over the next seven steps further improves prompts requiring fine-grained details such as color or texture, boosting performance on the Objects category in both ImageReward and BLIP-VQA. Additional gains come from the Closeness-Aware Concept Weight Modulation strategy, which consistently benefits all categories. We hypothesize that as the diffusion trajectory evolves, the latent state often drifts closer to one concept mode $c_{i}$; the modulation mechanism assigns $c_{i}$ a more negative weight, repelling the latent and mitigating concept dominance (Tunanyan et al., 2023) commonly observed in stable-diffusion.
BLIP-VQA ImageReward
Methods animals animals-objects objects animals animals-objects objects Avg.
SDXL 0.6951 0.8654 0.4926 0.7820 1.5474 0.6789 0.8435
+ resampling 0.7351 0.8666 0.4528 1.0881 1.6755 0.8263 0.9407
   + Corrector 0.7177 0.8794 0.4796 1.0630 1.6429 0.8949 0.9463
   + weight-modulation 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158 1.0118

### v2

# Steer Away from Mode Collisions: Improving Composition in Diffusion Models

@@ Abstract
We propose to improve multi‑concept prompt fidelity in text‑to‑image diffusion models. We begin with common failure cases—prompts like “a cat and a dog” that sometimes yields images where one concept is missing, faint, or colliding awkwardly with another. We hypothesize that this happens when the diffusion model drifts into mixed modes that over‑emphasize a single concept it learned strongly during training. Instead of re-training, we introduce a corrective sampling strategy that steers away from regions where the joint prompt behavior overlaps too strongly with any single concept in the prompt. The goal is to steer towards “pure” joint modes where all concepts can coexist with balanced visual presence. We further show that existing multi‑concept guidance schemes can operate in unstable weight regimes that amplify imbalance; we characterize favorable regions and adapt sampling to remain within them. Our approach, CO3, is plug‑and‑play, requires no model tuning, and complements standard classifier‑free guidance. Experiments on diverse multi-concept prompts indicate improvements in concept coverage, balance and robustness, with fewer dropped or distorted concepts compared to standard baselines and prior compositional methods. Results suggest that lightweight corrective guidance can substantially mitigate brittle semantic alignment behavior in modern diffusion systems. Code is available at https://github.com/debottam-dutta7/co3
@@ 1 Introduction
Recent diffusion models (Ho et al., 2020; Ramesh et al., 2022a; Rombach et al., 2022) have ushered significant breakthroughs in Text-to-Image (T2I) synthesis, producing high-fidelity images from textual descriptions. However, ensuring the generated images faithfully adhere to the prompt, a challenge known as semantic alignment (Chefer et al., 2023; taihang Hu et al., 2024; Liu et al., 2023), remains a problem. Concretely, for a given prompt $C$, T2I models like StableDiffusion (Rombach et al., 2022) sample from the modes (or high probability regions) of the learned distribution, $p(x\mid C)$. While such models can produce high resolution images in general, every so often, the results are surprisingly misaligned even for very simple prompts containing few concepts, e.g., CC=“a cat and a dog”. Diagnosing exactly why this behavior emerges periodically is difficult. It is conceivable that the complex training process in high dimensions, especially in conjunction with text embeddings, creates some problematic modes in $p(x\mid C)$.
We hypothesize that problematic modes in $p(x\mid C)$ arise when they overlap with modes of individual concept distribution $p(x\mid c_{i})$. Such an overlap biases the generation toward a single concept $c_{i}$, reducing the prominence of others. For instance, across images of $c_{1}=\text{``cat''}$ in the training dataset, a few may have an inconspicuous or partial $c_{2}=\text{``dog''}$ in the background. This image may still fall under the mode of $p(x\mid C)$. We attribute this to training instabilities and relatively less coverage of multi-concept prompts $C$, which cause the model to assign high probability even to weakly conforming images. Said differently, an image of a big cat and an inconspicuous dog can get assigned high probabilities under $p(x\mid C)$, causing semantic misalignment.
@@
Figure 1 illustrates the intuition behind our proposal. Our corrector distribution $\tilde{p}(x\mid C)$ assigns low probability to regions where $p(x\mid C)$ overlaps with individual $p(x\mid c_{i})$; we deem them as degenerate modes dominated by a single concept. By suppressing these overlaps, the corrector emphasizes pure $p(x\mid C)$ modes where all concepts coexist without one overwhelming the others. From a probabilistic perspective, this acts as a corrective factor: while $p(x\mid C)$ may assign high probability to weakly conforming images due to training noise or limited multi-concept data, dividing by the marginals removes this bias and sharpens the distribution toward genuine multi-concept samples. As a result, the modes we target are more semantically aligned and less prone to concept suppression or distortion.
Correction sampling from $\tilde{p}(x\mid C)$ can be achieved by composing scores from constituent component distributions $\nabla_{x}\log p(x\mid C),\{\nabla_{x}\log p(x\mid c_{i})\}_{i}$ (Liu et al., 2022). While there are many ways to compose, we analyse and show that composition through weighted sum of Tweedie-means—in the Tweedie denoised space—offers a more general framework that subsumes existing approaches.
In particular, we show that two classes of correction sampling—noise-resampling and latent correction (Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024; Kwon and Ye, 2025)—become special cases of Tweedie-mean composition under different weighting schemes. This allows us to design a hybrid composition framework that serves the purpose of resampling at time $T$, and then toggles to latent correction at later steps. By latent, we mean that the correction steps are accomplished in between each DDIM time step, allowing CO3(Concept Contrasting Corrector) to be a plug-and-play, model-agnostic and gradient-free approach for T2I models. Comparison of CO3 against SOTA baselines shows stronger semantic alignment to prompts (measured using multiple metrics), giving empirical evidence of our hypothesis.
Figure 1: The figure illustrates our hypothesis on mode overlap using a simple 2D toy example. (a) Two modes of the distribution $p(x\mid\texttt{"a cat and a dog"})$ (in green contour) has significant overlap with the modes of the individual concept distributions $p(x\mid\texttt{"a cat"})$ (in red contour) and $p_{t}(x\mid\texttt{"a dog"})$ (in orange contour). (b) The proposed (unnormalized) corrector distribution $p(x\mid\texttt{"a cat and a dog"})/(p(x\mid\texttt{"a cat"})p(x\mid\texttt{"a dog"}))$ suppresses these overlaps, steering the generation away from problematic modes. The arrows indicate the denoising directions.
@@ 2 Background
In practice, most T2I models adopt classifier-free guidance (CFG) (Ho and Salimans, 2022), where predictions from both the conditional and unconditional models are combined:
@@
where $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]:=x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}^{\lambda,c}$ is the Tweedie mean from the CFG noise at $t$. Thus, CFG in the Tweedie framework can be interpreted as modifying the denoised estimate $\hat{x}_{0}$ to $\hat{x}_{0}^{\lambda,c}$ before the deterministic step to $x_{t-1}$ (Chung et al., 2024; Kwon and Ye, 2023).
@@
$\tilde{p}_{0}(x_{0}\mid c_{1},\ldots,c_{K})\;\propto\;p(x_{0})\prod_{k=1}^{K}p_{0}(c_{k}\mid x_{0}).$ (9)
@@
3 CO3: Concept Contrasting Corrector
We aim to combine the strengths of correction-based approaches and composable diffusion. On one hand, correction-based methods are powerful and explicitly improve compositional alignment, but they are subject to the complex gradient of the base model. On the other hand, composable diffusion is fully model-agnostic, but suffers from poor performance because linear score composition is not consistent with the diffusion forward process Du et al. (2024).
To take advantage of both, assume that prompt $C$ can be decomposed into constituent concepts $\{c_{1},c_{2},\ldots,c_{K}\}$. We propose an explicit Concept Contrasting Corrector (CO3) distribution based on our hypothesis on mode overlap discussed in Sec. 1. Specifically, we define the corrector distribution at each timestep $t$ as:
@@ 3 CO3: Concept Contrasting Corrector
To sample from the unnormalized probability distribution in Eq. 12, an well-known approach is to compose the concept distributions $\{p_{t}(x_{t}\mid c_{k})\}_{k}$ in the space of score functions (Liu et al., 2022; Du et al., 2024). We break-away from this approach and compose the distribution in the Tweedie mean space; we show how this offers a more general framework for composition. In Tweedie-denoised space we define composition as:
@@
8:  $x_{t}\leftarrow\text{DDIM\_step}(x_{t})$ $\triangleright$ Reverse DDIM
@@
A valid compositional Tweedie mean $\tilde{x}_{tweedie}$ in Eq. 13 must be one that can be expressed in the definition given in equation 5 as DDIM goes through a series of Tweedie means for image generation as described in section 2. Proposition 1 shows that this is guaranteed only when the weights satisfy the normalization condition $\sum_{i}w_{i}=1$.
Proposition 1.
Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise ${\epsilon}_{t}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
@@
    CO3-corrector: $\tilde{\hat{x}}_{tweedie}$ can be expressed in the form of a tweedie-mean at time $t$, i.e. $\tilde{\hat{x}}_{tweedie}=x_{t}-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ if and only if $\sum_{k}w_{k}=1$. Here $\tilde{\lambda}=\lambda$ and CFG composed noise $\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$.
@@
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
@@ 4.1 Experimental Setup
Implementation details: We implement our method on SDXL (Podell et al., 2023) using a 50-step DDIM sampler with guidance scale $\lambda=5.0$. CO3 is applied during the first $20\%$ of denoising steps ($T_{c}=0.8T$), with the initial three steps reserved for resampling. For weight modulation, we set $\beta=0.8$. All experiments are performed on a single NVIDIA A6000 GPU (48GB). Additional implementation details are provided in the Appendix section A.4.
Evaluation benchmark and metrics: We first evaluate our method on simpler prompts (two concepts) using the Attend-Excite benchmark from Chefer et al. (2023) which contains categories of prompts: Animal–Animal, Animal–Object, and Object–Object. We then take the best performing baselines from this task and compare them with CO3 on more complex prompts (three or more concepts) from T2ICompbench Huang et al. (2023) and rare or novel concept prompts from RareBench (Park et al., 2024). Specifically, we consider ”Complex” category from T2ICompbench and ”Concat”, ”Relations” and ”Complex” from RareBench; the subset of categories which stress multi-concept interactions, attribute entanglement and rare-concept scenarios.
For evaluation, we adopt BLIP-VQA (Huang et al., 2023), ImageReward (Xu et al., 2023) and DINOv2 (Oquab et al., 2023) scores for simpler prompts. While BLIP-VQA and ImageReward (widely used in recent works (taihang Hu et al., 2024; Hu et al., 2024)) measure the faithfulness of generated-images to input-prompts, DINOv2 measures the quality of concepts generated by each model via a classification score. R2F Park et al. (2024) notes that BLIP-VQA can be unreliable in rare-concept scenarios; hence, following their recommendation, we report metrics that more closely reflect human judgement, namely, ImageReward and human evaluation (from 11 participants).
Comparison methods: We compare CO3 against state-of-the-art approaches, including optimization-based correction methods (Attend-Excite (Chefer et al., 2023), SynGen (Rassin et al., 2023), Divide-Bind (Liu et al., 2023), Magnet Zhuang et al. (2024), ToME (taihang Hu et al., 2024)); composable generation methods Comp-Diff (Liu et al., 2022), Tweediemix (Kwon and Ye, 2025); LLM augmented method R2F (Park et al., 2024) and the noise-optimization method InitNO (Guo et al., 2024). For fair comparison against Tweediemix’s sampling strategy, we implemented Tweediemix without its LORA-finetuned customization stage. All optimization-based approaches are architecture-dependent and require gradient computations at inference, with InitNO being model-agnostic but still incurring costly backpropagation. In contrast, only Composable Diffusion, Tweediemix and CO3 are purely sampling-based, model-agnostic, and completely gradient-free.
@@ 4.2 Experimental Results
Table 1: Quantitative comparison of different methods on two concept prompts from Chefer et al. (2023). We evaluate the generated images using three metrics: BLIP-VQA, ImageReward and DINOv2 scores on categories Animal-Animal(A-A), Animal-Object(A-O) and Object-Object(O-O). The top performing model is highlighted in and the 2nd best in .
Properties BLIP-VQA $\uparrow$ ImageReward $\uparrow$ DINOv2 $\uparrow$
train-free grad-free model-agnostic A-A A-O O-O A-A A-O O-O A-A A-O O-O
SD1.5(Rombach et al., 2022) - 0.3239 0.5958 0.2730 -0.2733 0.4262 -0.5521 0.2163 0.2651 0.3010
Attend-excite(Chefer et al., 2023) 0.6980 0.7865 0.5155 0.8244 1.2380 0.8741 0.2245 0.2599 0.2921
SynGenRassin et al. (2023) 0.3348 0.7689 0.6595 -0.2445 0.9838 0.7315 0.2187 0.2606 0.2790
Divide & Bind 0.7201 0.8399 0.5887 0.8499 1.2516 0.8134 0.2029 0.2481 0.2865
InitNO(Guo et al., 2024) 0.7264 0.7998 0.5406 1.0082 1.3927 1.1383 0.2252 0.2654 0.2953
SDXLPodell et al. (2023) - 0.6950 0.8654 0.4926 0.7820 1.5574 0.6789 0.2234 0.2618 0.2629
Magnet Zhuang et al. (2024) 0.6782 0.8744 0.5398 0.6450 1.5895 0.7856 0.2122 0.2636 0.2624
SynGen(SDXL)Rassin et al. (2023) 0.6816 0.8578 0.4652 0.6998 1.5622 0.6441 0.2221 0.2650 0.2710
ToMe(taihang Hu et al., 2024) 0.6257 0.8808 0.6440 0.3895 1.5736 1.0118 0.2226 0.2620 0.2545
TweediemixKwon and Ye (2025)111adapted without the (LORA finetuned) concept-customization stage. 0.7390 0.8059 0.4683 1.0023 1.3127 0.7959 0.2519 0.3105 0.3237
Comp-diff(Liu et al., 2022) 0.2846 0.5656 0.4529 -1.1399 -0.2068 -0.0955 0.2497 0.2756 0.2908
CO3(Ours) 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158 0.2623 0.3259 0.3283
Table 2: Quantitative results on multi-concept prompts from T2ICompbench and rare-concepts from RareBench. We report ImageReward (IR) and Human Evaluation (HE) scores.
T2ICompbench RareBench
Complex Concat Relations Complex
Method IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$
SDXL (Podell et al., 2023) 0.3083 0.5983 0.1018 0.5416 0.3607 0.5016 1.1907 0.6283
SynGen(SDXL) (Rassin et al., 2023) 0.2889 0.5133 0.0576 0.3177 1.1452
ToMe (taihang Hu et al., 2024) 0.2891 0.5866 -0.2041 0.4500 -0.0632 0.4983 1.0200 0.48666
R2F (Park et al., 2024) 0.3561 0.4816 -0.0162 0.4300 0.3620 0.5200 1.1542 0.53666
CO3(Ours) 0.4406 0.8033 0.1610 0.6167 0.2588 0.6200 1.2211 0.61166
@@
Simpler prompts (Attend-Excite): Table 1 demonstrates the extent to which CO3, despite being model-agnostic and optimization-free, competes or outperforms optimization-based and model-agnostic baselines on ImageReward. Gains are substantive in the Animal–Animal and Animal–Object categories. On BLIP-VQA, CO3 delivers the best performance in these two categories and improves over the base SDXL on the Object category, which contains prompts of the form: “a [color1] [object1] and a [color2] [object2].” Unlike methods such as ToMe taihang Hu et al. (2024), SynGen (Rassin et al., 2023), Attend-Excite (Chefer et al., 2023), and Divide-Bind Liu et al. (2023), which rely on model-specific architectures or explicit attention-based binding losses, CO3 is architecture-independent, sampling-based, and free of explicit subject–attribute binding. Interestingly, by simply targeting “pure” modes under $p(x\mid C)$, CO3 discovers layouts that better capture multi-object relations (Animals–Animals, Animal–Objects) and finds high-probability regions that preserve subject–attribute bindings (Objects).
@@
Multiple and rare-concept prompts (T2ICompbench and RareBench): Table 2 shows that CO3 consistently outperforms or matches the best-performing baselines across all categories and metrics. We note that CO3 is not designed specifically for rare-concept scenarios—it is a general compositional sampling correction that does not rely on rarity-aware priors or LLM-based prompt augmentation. Yet, CO3 matches or exceeds R2F on RareBench, despite R2F being tailored for rare-concept generation. We believe this indicates that the early-alignment phase of CO3, which mitigates concept dominance and encourages sampling from “pure” modes, is naturally beneficial for rare-concept generation as well.
These performance gains empirically validates our hypothesis on mode overlap and also highlights CO3’s effectiveness in correcting the compositional limitations of the base model.
Figure 4: Qualitative comparison of CO3 with competing methods on complex and rare concept prompts.